coldz16:

Given the trinomial, what is the value of the coefficient 'B' in the factored form 2x2 + 14xy − 36y2 = 2(x + By)(x − 2y)

jhonyy9:

2x^2 +14xy -36y^2 = 2(x+By)(x-2y) 2(x^2 +7xy -18y^2) = 2(x+By)(x-2y) divide both sides by 2 and will get x^2 +7xy -18y^2 = (x+By)(x-2y) so than you check when you multiply the By by (-2y) you need getting -18y^2 so By * (-2y) = -18y^2 how you think when will get this needed -18y^2 than what number is in the place of B ? hope helped understandably